OpenStudy (anonymous):

What is the area of a regular pentagon with a side length of 9.4 feet and an apothem length of 6.5 feet?

OpenStudy (anonymous):

 From the drawing you get five equal triangles with base 9.4 feet, and height 6.5 feet. \[A_{pentagon}=5*A_{Triangle}\] \[A_{pentagon}=5*((b*h)/2)\]

OpenStudy (anonymous):

thanks (: it really did